The nursing staff is currently composed of 20 female and 5 male nurses. over the next year the clinic hopes to increase the male nursing staff to equal at least 25% of the total nursing staff. if everyone now working retains his or her job what is the least number of people the clinic will need to hire?
Answer #1

There are 20 females and 5 males so,

Total staff = 20 + 5 = 25

There is 25% increment in people of total staff in next year.

So, for the next year number of people will be increased by,

25 * 0.25 = 6.25

Approximately 7 people

Hence, next year clinic will need to hire at least 7 people

Note: the key here is that everyone is retaining his or her job so you are adding to the total. you can do this by guessing and checking right now there are 5/25 males or 20 percent of the nursing staff. if you add just one more the total will be 6/25 male nurses or 23 percent adding another brings you to 7/27 male nurses or just more than 25 percent of the nursing staff so adding 2 would solve the diversity